# HAIFIT: Human-Centered AI for Fashion Image Translation

### v2

@@ Abstract
In the realm of fashion design, sketches serve as the canvas for expressing an artist’s distinctive drawing style and creative vision, capturing intricate details like stroke variations and texture nuances. The advent of sketch-to-image cross-modal translation technology has notably aided designers. However, existing methods often compromise these sketch details during image generation, resulting in images that deviate from the designer’s intended concept. This limitation hampers the ability to offer designers a precise preview of the final output. To overcome this challenge, we introduce HAIFIT, a novel approach that transforms sketches into high-fidelity, lifelike clothing images by integrating multi-scale features and capturing extensive feature map dependencies from diverse perspectives. Through extensive qualitative and quantitative evaluations conducted on our self-collected dataset, our method demonstrates superior performance compared to existing methods in generating photorealistic clothing images. Our method excels in preserving the distinctive style and intricate details essential for fashion design applications.  [^1] Code is available at https://github.com/ExponentiAI/HAIFIT/.
@@ I Introduction
To address these challenges, we introduce HAIFIT (Human-Centered AI for Fashion Image Translation), a GAN-based creative generation network designed for sketch-to-image translation in fashion design. HAIFIT aims to empower designers to foster inspiration by preserving as much detail as possible from their sketches in the generated images. Specifically, HAIFIT focuses on converting clothing sketches into high-quality realistic clothing images using a pyramidal generative adversarial network. To facilitate this conversion, we have developed a customized Multi-scale Feature Fusion Encoder (MFFE) that maps the reference hand-drawn sketches and generated images into a shared latent space. MFFE comprises a Shallow Convolution Module (SCM) and an Abstract Feature Representation Module (AFRM), which enables the capture of long-term dependencies between different regions and improves the global correlation in feature representations. Furthermore, we enhance the fine-grained features among generated images of various scales by integrating the Cross-level Skip Connection Module (CSCM) within the feature space across different layers. By employing multi-scale discriminators, we ensure the preservation of global structural integrity while emphasizing fine details, resulting in more realistic generated images. The key contributions of our work are summarised as follows:
  • Methodology. To the best of our knowledge, we are the pioneers in treating sketch features as sequence inputs, ensuring consistency by inputting features from both forward and reverse feature sequences. This innovative approach enhances the representation of details in the generated images.
@@ III Methodology
Algorithm 1 The PGGAN Training Process.
0:  $n$: Number of multi-layers
0:  $G=\{G_{1},G_{2},...,G_{n}\}$: Generator networks
0:  $D=\{D_{1},D_{2},...,D_{n}\}$: Discriminator networks
0:  $I=\{I_{1},I_{2},...,I_{n}\}$: Image resolutions
0:  $k$: Number of training iterations
1:  $G\leftarrow G_{1}$, $D\leftarrow D_{1}$, $I\leftarrow I_{1}$
2:  Initialize $G_{1}$ and $D_{1}$ with random weights
3:  for $i=1$ to $n$ do
4:     Training $D_{i}$ for $k$ iterations with fixed $G_{i}$
5:     Training $G_{i}$ for $k$ iterations with fixed $D_{i}$
6:     if $I_{i}<I_{n}$ then
[…]
8:        Initialize new layers $G_{i+1}$ and $D_{i+1}$ for higher resolution
9:        Update weights of existing layers in $G$ and $D$
10:     end if
11:  end for
@@
In this study, we present HAIFIT, as depicted in Fig. 2, which is based on PGGAN [13] and comprises two core components: the Multi-scale Feature Fusion Encoder (MFFE), and the Cross-level Skip Connection Module (CSCM). The MFFE is used to get global contour features and abstract intent features. Then the feature maps are resized and concatenated at multiple scales before being fed into the multi-scale pyramid generator, which can produce realistic images with complete structures and exquisite details.
III-A Preliminaries
The traditional GANs [6] enhance the model’s generation capabilities by adversarially training the generator and discriminator. However, they still face significant challenges in terms of image generation quality and training process stability. To address these issues, PGGAN [13] introduces a progressive growth training method. This approach starts from low-resolution images and gradually increases the number of layers and image resolution of both the generator and discriminator. This progressive growth training process not only simplifies the model’s convergence to generate higher-quality and more flexible-resolution images, but also mitigates issues such as mode collapse and gradient disappearance during the training process. Consequently, we utilize the PGGAN as the framework for our approach.
The specific algorithmic process is outlined in Algorithm 1. We commence training the initial generator $G_{1}$ and initial discriminator $D_{1}$ from the initial resolution $I_{1}$, gradually increasing the resolution and continuing training by adding new $G_{i+1}$ and $D_{i+1}$ layers. This process continues until the final resolution $G_{n}$ and $D_{n}$ model training is completed, ultimately yielding the desired generated image.
Figure 2: The illustration of our HAIFIT architecture. Given $M$ paired sketch-image samples $\{(x_{i},y_{i})\in(X,Y)\}^{M}_{i=1}$, our goal is to map images from hand-drawn sketches domain $X$ to realistic images domain $Y$, while preserving as many line details from the sketch $X$ as possible in the generated image $\overline{Y}$.
III-B Multi-scale Feature Fusion Encoder
@@
The Multi-scale Feature Fusion Encoder (MFFE) consists of two modules: the Shallow Convolutional Module (SCM) and the Abstract Feature Representation Module (AFRM). As illustrated in Fig. 3, the SCM employs a convolutional network to learn essential global contour information, such as the structure of the object. Simultaneously, inspired by the LSTM architecture  [10], we design the AFRM to learn long-range dependencies in key areas of the input sketch from various perspectives, which can effectively capture style features and geometric features of strokes. By combining these features, we can effectively capture the consistency of objects’ complete structure and intricate textures within the latent space.
@@
As shown in Fig. 2, in the bottom layer, only the sketch features are inputted. In the other layers, we perform element-wise addition between the generated image features and the sketch features, and then input the resulting summed features into the MFFE. In the SCM we employ a sequence of convolutional layers, denoted as $Con$, to extract the global contour features and shape information $\widetilde{x}$ from the input sketch $x_{s}$. Here, $x_{s}\in\mathcal{R}^{b\times 3\times m\times m}$, $\widetilde{x}\in\mathcal{R}^{b\times 256\times m/4\times m/4}$, $b$ is the batch size, and $m$ depends on the network layer, so we can express this process as $\widetilde{x}=Con(x_{s})$. Specifically, the SCM extracts features in a blockwise manner, which will disrupt the continuity of strokes and result in pixel-level contour features. Since sketches are continuous and composed of multiple lines in different directions, it is crucial to consider the correlation between adjacent lines in order to effectively learn the line features of complex sketches.
@@
The MFFE is employed to learn features from different perspectives of the sketches, which enables us to capture variations in lines and abstract painting intentions in crucial areas of the sketches. In the task of sketch-to-image generation, it is vital for the encoder to acquire diverse line patterns and painterly intent in clothing patterns and details. Failure to do so can severely impact the visual quality of the synthesized image, even if the overall generation quality is satisfactory. As a result, to further ensure the consistency of the generated images, we introduce the concept of reverse sketch feature input. This is because a sketch should yield consistent results regardless of the direction it is interpreted from. Therefore, the aforementioned process can be summarized as follows:
$\begin{split}\overline{x}_{dir}^{\prime}&=Cat(LSTM[\widetilde{x}_{0},\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3}])\\ \overline{x}_{rev}^{\prime}&=Cat(LSTM[\widetilde{x}_{0},\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3}])\\ \overline{x}&=Conv_{1\times 1}(Cat(\overline{x}_{dir}^{\prime},\overline{x}_{rev}^{\prime}))\end{split}$ (1)
@@
Figure 3: An overview of the Multi-scale Feature Fusion Encoder (MFFE).
III-C Cross-level Skip Connection Module
@@
$Y_{n}=\begin{cases}G_{n}(X_{n}),&\mbox{if {n} = 1}\\ G_{n}(X_{n},Y_{n-1}),&\mbox{if {n} > 1}\end{cases}$ (3)
The generator $G_{n}$ consists of a residual block and multiple convolutional layers. In the coarsest layer, the residual block comprises four convolutional layers. In addition, in the pyramid network, as the number of pixels increases from bottom to top, the residual block progressively adds one convolutional layer at each level. This progressive increment facilitates the learning of sketch feature representations. Through numerous experiments, it has been observed that incorporating the CSCM between different layers in the pyramid network can enhance the quality of the generated images, which is illustrated in Fig. 4. In detail, We extract the output feature ${X^{\prime}}_{n-1}$ from the last layer of the residual block in generator $G_{n-1}$ and concatenate it with the output feature ${X}_{n}$ from the $n$-th layer of the MFFE. Subsequently, we use $\overline{X}_{n}$ as the input for generator $G_{n}$, where $\overline{X}_{n}\in\mathcal{R}^{b\times 256\times m\times m}$. In principle, the multi-level features of the same synthetic image should exhibit similarity. By employing cross-level feature operations, we can extract similar details from the features generated by different level generators and incorporate them as attention maps into the new feature map $X_{n}$. This facilitates image translation, resulting in more detailed and high-quality images.
@@
Figure 4: An overview of the Cross-level Skip Connection Module (CSCM).
III-D Loss Function
@@
IV Experiments
IV-A HAIFashion Dataset
@@
Common publicly available fashion datasets like FashionGEN [23] and DeepFashion [19] predominantly feature realistic fashion images sourced directly from real-world photos. These images often include various backgrounds and individuals wearing the clothing in social settings. However, images with backgrounds pose significant challenges in generating single clothing sketches, which may not meet our requirements. The comparison between these datasets is illustrated in Fig. 5. In this study, we introduce a self-collected clothing dataset named HAIFashion. HAIFashion comprises 3100 pairs of fashion clothing image-sketch datasets. We achieved this by crawling 3100 background-free fashion clothing images from the internet and employing the algorithm [31] to generate designer-style clothing sketches, facilitating our experiments.
IV-B Implementation Details
We implement our model on Pytorch and all experiments are performed on a single NVIDIA GeForce RTX4090 GPU. We utilize the Adam optimizer with a learning rate of 1e-4 for the generator and 5e-4 for the discriminator, the batch size is set to 8. The learning rates decayed at the halfway point every 100 epochs. The training process continued until the SSIM [30] showed no improvement for 10 iterations. For the dataset, we use our proposed HAIFashion (§ IV-A) and divide it into 2500 (600) for model training and testing respectively.

### v3

@@ Abstract
In the realm of fashion design, sketches serve as the canvas for expressing an artist’s distinctive drawing style and creative vision, capturing intricate details like stroke variations and texture nuances. The advent of sketch-to-image cross-modal translation technology has notably aided designers. However, existing methods often compromise these sketch details during image generation, resulting in images that deviate from the designer’s intended concept. This limitation hampers the ability to offer designers a precise preview of the final output. To overcome this challenge, we introduce HAIFIT, a novel approach that transforms sketches into high-fidelity, lifelike clothing images by integrating multi-scale features and capturing extensive feature map dependencies from diverse perspectives. Through extensive qualitative and quantitative evaluations conducted on our self-collected dataset, our method demonstrates superior performance compared to existing methods in generating photorealistic clothing images. Our method excels in preserving the distinctive style and intricate details essential for fashion design applications. In addition, our method also has obvious advantages in model training and inference speed, contributing to reducing designers’ time costs and improving design efficiency. Code is available at https://github.com/ExponentiAI/HAIFIT/.
@@ I Introduction
To address these challenges, we introduce HAIFIT (Human-Centered AI for Fashion Image Translation), a GAN-based creative generation network designed for sketch-to-image translation in fashion design. HAIFIT aims to empower designers to foster inspiration by preserving as much detail as possible from their sketches in the generated images. Specifically, HAIFIT focuses on converting clothing sketches into high-quality realistic clothing images using a pyramidal generative adversarial network. To facilitate this conversion, we have developed a customized Multi-scale Feature Fusion Encoder (MFFE) that maps the reference hand-drawn sketches and generated images into a shared latent space. MFFE comprises a Shallow Convolution Module (SCM) and an Abstract Feature Representation Module (AFRM), which enables the capture of long-term dependencies between different regions and improves the global correlation in feature representations. Furthermore, we enhance the fine-grained features among generated images of various scales by integrating the Cross-level Skip Connection Module (CSCM) within the feature space across different layers. By employing multi-scale discriminators, we ensure the preservation of global structural integrity while emphasizing fine details, resulting in more realistic generated images.
The key contributions of our work are summarised as follows:
  • Methodology. To the best of our knowledge, we are the pioneers in treating sketch features as sequence inputs, ensuring consistency by inputting features from both forward and reverse feature sequences. This innovative approach enhances the representation of details in the generated images. Moreover, we are the first to propose a Pyramid Diffusion Model, which uses a progressive generation approach to improve the quality of model generation.
@@
The rest of the article is structured as follows: Sec. II provides an overview of the work and technological progress related to image translation. Sec. III presents the technical details of our approach. Sec. IV introduces a version of our method extended to a diffusion model. Sec. V presents the experimental results of our approach. Sec. VI summarizes our work.
Figure 2: (i) [Left] The illustration of our HAIFIT architecture. Given $M$ paired sketch-image samples $\{(x_{i},y_{i})\in(X,Y)\}^{M}_{i=1}$, our goal is to map images from hand-drawn sketches domain $X$ to realistic images domain $Y$, while preserving as many line details from the sketch $X$ as possible in the generated image $\overline{Y}$. (ii) [Top Right] An overview of the Multi-scale Feature Fusion Encoder (MFFE). (iii) [Bottom Right] An overview of the Cross-level Skip Connection Module (CSCM).
@@ II-B Generative Adversarial Networks
II-C Diffusion Models
Different from GANs, the diffusion model [10] realizes image generation from a novel perspective. It comprises two stages: the forward noise-adding stage and the reverse denoising stage. In the forward noise-adding stage, Gaussian noise is continuously added to the initial input data to make it approximately Gaussian. In the reverse denoising stage, the model is tasked with recovering the original input data by learning to gradually reverse the denoising process. Diffusion models are widely appreciated for the quality and diversity of the samples they generate, but they also impose a severe computational burden and time cost. DDIM [30] addresses the issue of single-step sampling in the inference process through skip sampling. Subsequent advancements like ADM [5] further demonstrate the superiority of the diffusion model over traditional GAN methods in image synthesis. Another notable approach by Ho et al. [11] proposes a diffusion model training method without classifier guidance. Palette [28] introduces a unified framework for image-to-image translation based on the conditional diffusion model, achieving superior results across multiple image generation subtasks. However, they still encounter significant model training costs. LDM [26] addresses this by significantly reducing model training costs and opening up new avenues for subsequent research by placing the diffusion process in latent space. DiT [24] explores a new class of diffusion models based on transformer architecture, enhancing the scalability and quality of diffusion models. Overall, these studies continuously advance the field of image generation and offer new possibilities for natural and realistic image synthesis.
@@ III Methodology
In this study, we present HAIFIT, as depicted in Fig. 2, which draws on the training strategy of PGGAN [16] and comprises two core components: the Multi-scale Feature Fusion Encoder (MFFE), and the Cross-level Skip Connection Module (CSCM). The MFFE is used to get global contour features and abstract intent features. Then the feature maps are resized and concatenated at multiple scales before being fed into the multi-scale pyramid generator, which can produce realistic images with complete structures and exquisite details.
III-A Multi-scale Feature Fusion Encoder
@@
The Multi-scale Feature Fusion Encoder (MFFE) consists of two modules: the Shallow Convolutional Module (SCM) and the Abstract Feature Representation Module (AFRM). As illustrated in Fig. 2 (ii), the SCM employs a convolutional network to learn essential global contour information, such as the structure of the object. Simultaneously, inspired by the LSTM architecture  [12], we design the AFRM to learn long-range dependencies in key areas of the input sketch from various perspectives, which can effectively capture style features and geometric features of strokes. By combining these features, we can effectively capture the consistency of objects’ complete structure and intricate textures within the latent space.
@@
As shown in Fig. 2 (i), in the bottom layer, only the sketch features are inputted. In the other layers, we perform element-wise addition between the generated image features and the sketch features, and then input the resulting summed features into the MFFE. In the SCM we employ a sequence of convolutional layers, denoted as $Con$, to extract the global contour features and shape information $\widetilde{x}$ from the input sketch $x_{s}$. Here, $x_{s}\in\mathcal{R}^{b\times 3\times m\times m}$, $\widetilde{x}\in\mathcal{R}^{b\times 256\times m/4\times m/4}$, $b$ is the batch size, and $m$ depends on the network layer, so we can express this process as $\widetilde{x}=Con(x_{s})$. Specifically, the SCM extracts features in a blockwise manner, which will disrupt the continuity of strokes and result in pixel-level contour features. Since sketches are continuous and composed of multiple lines in different directions, it is crucial to consider the correlation between adjacent lines in order to effectively learn the line features of complex sketches.
@@
The MFFE is employed to learn features from different perspectives of the sketches, which enables us to capture variations in lines and abstract painting intentions in crucial areas of the sketches. In the task of sketch-to-image generation, it is vital for the encoder to acquire diverse line patterns and painterly intent in clothing patterns and details. Failure to do so can severely impact the visual quality of the synthesized image, even if the overall generation quality is satisfactory. As a result, to further ensure the consistency of the generated images, we introduce the concept of reverse sketch feature input. This is because a sketch should yield consistent results regardless of the direction it is interpreted from. As shown in Fig. 2, we provide the structural details related to the AFRM in MFFE. We serialize the feature $\widetilde{x}$ obtained from the SCM and input it into the AFRM through forward and reverse feature inputs. The output features $\overline{x}_{for}^{\prime}$ and $\overline{x}_{rev}^{\prime}$ are then fused to obtain the unified sketch abstract stroke feature $\overline{x}$, which is subsequently used in the later stages of the model. Therefore, the aforementioned process can be summarized as follows:
$\begin{split}\overline{x}_{for}^{\prime}&=Cat(LSTM[\widetilde{x}_{0},\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3}])\\ \overline{x}_{rev}^{\prime}&=Cat(LSTM[\widetilde{x}_{3},\widetilde{x}_{2},\widetilde{x}_{1},\widetilde{x}_{0}])\\ \overline{x}&=Conv_{1\times 1}(Cat(\overline{x}_{for}^{\prime},\overline{x}_{rev}^{\prime}))\end{split}$ (1)
@@
III-B Cross-level Skip Connection Module
@@
$Y_{n}=\begin{cases}G_{n}(X_{n}),&\mbox{if {n} = 1}\\ G_{n}(X_{n},Y_{n-1}),&\mbox{if {n} \textgreater 1}\end{cases}$ (3)
The generator $G_{n}$ consists of a residual block and multiple convolutional layers. In the coarsest layer, the residual block comprises four convolutional layers. In addition, in the pyramid network, as the number of pixels increases from bottom to top, the residual block progressively adds one convolutional layer at each level. This progressive increment facilitates the learning of sketch feature representations. Through numerous experiments, it has been observed that incorporating the CSCM between different layers in the pyramid network can enhance the quality of the generated images, which is illustrated in Fig. 2 (iii). In detail, We extract the output feature ${X^{\prime}}_{n-1}$ from the last layer of the residual block in generator $G_{n-1}$ and concatenate it with the output feature ${X}_{n}$ from the $n$-th layer of the MFFE. Subsequently, we use $\overline{X}_{n}$ as the input for generator $G_{n}$, where $\overline{X}_{n}\in\mathcal{R}^{b\times 256\times m\times m}$. In principle, the multi-level features of the same synthetic image should exhibit similarity. By employing cross-level feature operations, we can extract similar details from the features generated by different level generators and incorporate them as attention maps into the new feature map $X_{n}$. This facilitates image translation, resulting in more detailed and high-quality images.
@@
Algorithm 1 The Diffusion HAIFIT Training Process.
0:  $S$: Input Sketches
0:  $E$: Sketch Encoder network
0:  $\epsilon$: Gaussian Noise
0:  $step$: Model training steps
0:  $t\in[0,1000]$: Diffusion moment
0:  $G_{i}\in\{G_{1},G_{2},G_{3},G_{4}\}$: Generator networks
0:  $\bar{Y}_{i}\in\{\bar{Y}_{1},\bar{Y}_{2},\bar{Y}_{3},\bar{Y}_{4}\}$: Generated Images
1:  Randomly initialize $\epsilon$ as Input
2:  for $i=1$ to 4 do
3:     Utilize $E$ to encode $S$ as supervised conditions
4:     for 0 to $step$ do
5:        Randomly initialize $t$
6:        Train $G_{i}$ to predict the $t$-th moment noise $\bar{\epsilon}_{t}$
7:        Compute loss between $\bar{\epsilon}_{t}$ and real noise $\epsilon_{t}$
8:        Update parameters of $G_{i}$ using gradient descent
9:     end for
10:     if $i<4$ then
11:        Freeze parameters in $G_{i}$ and use it for inference
12:        Upsample $\bar{Y}_{i}$ to double its original resolution and concatenate with new resolution $\epsilon$ as Input
13:     end if
14:  end for
15:  Utilize all $\{G_{i}\}_{i=1}^{4}$ for inference generation
III-C Loss Function
@@
IV Pyramid Diffusion Model
In recent years, the Diffusion models [10, 30, 26] have garnered significant attention from researchers due to their potent generation capabilities, leading to rapid advancements and widespread adoption across various fields. Building upon this, we extend our GAN-based HAIFIT model to incorporate a diffusion model version: the Pyramid Diffusion Model.
Figure 3: The illustration of our HAIFIT with Diffusion model architecture.
IV-A Preliminaries of DDPM
The Denoising Diffusion Probabilistic Model (DDPM) comprises a forward process and a reverse process. In the forward process, noise $\epsilon$ is iteratively added to the original image $X_{0}$, resulting in a noisy image $X_{T}$. The noise $\epsilon$ added at each time step $t$ is sampled from a Gaussian distribution $\mathcal{N}(0,I)$. When $T$ is sufficiently large, we believe that $X_{T}$ approximately conforms to the Gaussian distribution, here we set $T=1000$. This transmission can be mathematically represented as $q(X_{t}|X_{t-1})=\mathcal{N}(\sqrt{1-\beta_{t}}X_{t-1},\beta_{t}I)$, $\{\beta_{t}\}_{t=0}^{T}\in(0,1)$. Therefore, the noised image $X_{t}$ at $t$-th moment in the forward process can be expressed as:
$X_{t}=\sqrt{\bar{\alpha}_{t}}X_{0}+\sqrt{(1-\bar{\alpha}_{t})}\epsilon,\epsilon\sim\mathcal{N}(0,I)$ (10)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$.
In the reverse process stage, the noisy image $X_{T}$ is continuously denoised to revert it back to the original input image $X_{0}$. This is achieved by training a model $\theta$ to predict the noise $\epsilon$ added to at each time step $t$, thereby removing the corresponding noise $\epsilon$. Therefore, the approximate probability distribution of the image $X_{t-1}$ at time $t$ can be expressed as:
$p_{\theta}(X_{t-1}|X_{t})=\mathcal{N}(X_{t-1};\mu_{\theta}(X_{t},t),\sigma^{2}I)$ (11)
To summarize, the optimization objective function of the diffusion model training process is:
$\mathcal{L}(\theta)=\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}\Big{]}$ (12)
IV-B Diffusion HAIFIT
As presented in Fig. 3, We utilize the diffusion method to hierarchically train each layer in HAIFIT. Beginning from the $L_{1}$ layer, we progressively train upwards to the $L_{4}$ layer. Upon completing the training of the previous layer network, we freeze the parameters of the network and add a new layer network upwards for further training until all layers of the model are trained. Specifically, we initialize random noise $\epsilon$ and then train $G_{1}$ to predict the noise $\epsilon_{t}$ at each moment $t$ and perform denoising operations on $X_{t}$ until $G_{1}$ effectively simulates the noise value added at each moment $t$, where $t\in[0,1000]$. During this process, we employ an encoder $E$ to encode the features of the corresponding sketch $S$ for supervised diffusion model training. This encoder is a backbone based on a pre-trained ResNet [8]. We adjust the dimension of the final encoded feature output by adding a trainable linear layer behind the backbone. Once $G_{1}$ training is complete, we freeze the parameters in $G_{1}$ and utilize it for inference, subsequently training the parameters in $G_{2}$. The input of $G_{2}$ contains both the noised image $X_{t}$ at the resolution of this layer and the low-resolution image $\bar{Y}_{1}$ generated by $G_{1}$. We upsample $\bar{Y}_{1}$ to double its original resolution and then concatenate it with the noised image $X_{t}$ in the channel dimension before inputting it into $G_{2}$. Following a similar procedure, we train the generators $G_{3}$ and $G_{4}$ successively. Upon completion of training, we utilize all layers in the model for inference generation. The principle of our Diffusion version of HAIFIT can be referenced in Algorithm 1.
V Experiments
V-A HAIFashion Dataset
@@
Common publicly available fashion datasets like FashionGEN [27] and DeepFashion [22] predominantly feature realistic fashion images sourced directly from real-world photos. These images often include various backgrounds and individuals wearing the clothing in social settings. However, images with backgrounds pose significant challenges in generating single clothing sketches, which may not meet our requirements. The comparison between these datasets is illustrated in Fig. 4. In this study, we introduce a self-collected clothing dataset named HAIFashion. HAIFashion comprises 3100 pairs of fashion clothing image-sketch datasets. We achieved this by crawling 3100 background-free fashion clothing images from the internet and employing the sketch generate algorithm [35] and the paired sketch-image dataset [15] to generate designer-style clothing sketches, facilitating our experiments.
V-B Implementation Details
We implement our model on Pytorch and all experiments are performed on a single NVIDIA GeForce RTX4090 GPU. We utilize the Adam optimizer with a learning rate of 1e-4 for the generator and 5e-4 for the discriminator, the batch size is set to 8. The learning rates decayed at the halfway point every 100 epochs. The training process continued until the SSIM [34] showed no improvement for 10 iterations. For the dataset, we use our proposed HAIFashion (§ V-A) and divide it into 2500 (600) for model training and testing respectively.